Can you please put into detail of what symbols would need to go into the calculator?
Activity Two: Discrete Probability Distributions in Action Example #2) Suppose a health insurance company can resolve 60% of claims usinga computerised system, the remaining needing work by humans. day, 10 claims arrived, assuming claims are independent, what is the probablity that (Adapted from NewBold et al., 2003, p. 149): On a particular Q2.1) Either 3 or 4 (inclusive) claims require work by a human? Q2.2) No more than 9 claims require work by a human? Present your answers as a percentage to four decimal place s.
Answer #1

Ans:

Use Binomial distribution with n=10 and p=1-0.6=0.4

P(x=k)=10Ck*0.4k*0.610-k

1)

P(x=3 or 4)=P(x=3)+P(x=4)

=10C3*0.43*0.67+10C4*0.44*0.66

=0.2150+0.2508

=0.4658

2)

P(no more than 9)=P(x<=9)=0.9999